### v2

# Scalable Spatio-Temporal SE(3) Diffusion for Long-Horizon Protein Dynamics

Abstract
Molecular dynamics (MD) simulations remain the gold standard for studying protein dynamics, but their computational cost limits access to biologically relevant timescales. Recent generative models have shown promise in accelerating simulations, yet they struggle with long-horizon generation due to architectural constraints, error accumulation, and inadequate modeling of spatio-temporal dynamics. We present STAR-MD (Spatio-Temporal Autoregressive Rollout for Molecular Dynamics), a scalable -equivariant diffusion model that generates physically plausible protein trajectories over microsecond timescales. Our key innovation is a causal diffusion transformer with joint spatio-temporal attention that efficiently captures complex space-time dependencies while avoiding the memory bottlenecks of existing methods. On the standard ATLAS benchmark, STAR-MD achieves state-of-the-art performance across all metrics–substantially improving conformational coverage, structural validity, and dynamic fidelity compared to previous methods. STAR-MD successfully extrapolates to generate stable microsecond-scale trajectories where baseline methods fail catastrophically, maintaining high structural quality throughout the extended rollout. Our comprehensive evaluation reveals severe limitations in current models for long-horizon generation, while demonstrating that STAR-MD’s joint spatio-temporal modeling enables robust dynamics simulation at biologically relevant timescales, paving the way for accelerated exploration of protein function.
[…]
1 Introduction
Protein functions emerge from conformational dynamics – the continuous structural changes that underlie important biological processes such as catalysis, binding, and allosteric regulations [mccammon1984proteindynamics, berendsen2000collectivedynamics]. Classical molecular dynamics (MD) simulation remains the gold standard and relies on physical models to integrate atomic motions over time using Newtonian mechanics. However, the need for small integration steps ( femtoseconds) severely limits its practicality for exploring the microsecond–millisecond timescales often required to capture biologically relevant events. Recent advances in protein structure prediction [jumper2021_af2, abramson2024accurate, lin2023evolutionary] and generative modeling for conformational dynamics [jing2024generative_mdgen, cheng20254d, shen2025simultaneous_confrover, costa2024equijump] offer promising data-driven approaches to accelerate simulations by learning the dynamics directly from data and generating trajectories at coarser temporal resolutions. However, existing methods are still constrained by short time horizons (typically up to nanoseconds) and struggle to scale to larger proteins. These limitations highlight an urgent need for generative models capable of producing physically plausible protein trajectories over extended timescales while remaining computationally efficient.
Scaling protein dynamics modeling to long timescales requires both an efficient structural representation and an architecture capable of capturing complex spatio-temporal dependencies. Current approaches often rely on pairwise residue representations or computationally expensive architectures (e.g., AlphaFold2-style triangular attention), leading to quadratic memory growth and cubic computational cost with respect to protein size. These challenges become even more pronounced when accounting for spatio-temporal coupling across multiple conformations during trajectory modeling. As a result, most existing models treat spatial and temporal components separately, employing interleaved “spatial” and “temporal” modules [jing2024generative_mdgen, cheng20254d, shen2025simultaneous_confrover], which limits their expressiveness in capturing coupled dynamics. These limitations confine current models to small proteins and short simulation horizons, ultimately hindering their ability to learn long-range temporal dependencies and generate high-quality conformations over extended rollouts.
We introduce Spatio-Temporal Autoregressive Rollout for Molecular Dynamics (STAR-MD), an -equivariant autoregressive diffusion model for generating physically plausible protein trajectories over microsecond timescales, even for large protein systems. At its core, STAR-MD employs a causal diffusion transformer with joint spatio-temporal attention, enabling improved autoregressive generation by dynamically computing attention over historical context during diffusion and capturing long-range dependencies through expressive yet memory-efficient spatio-temporal modeling.
Our main contributions are summarized as follows:
We present STAR-MD, a novel autoregressive diffusion transformer that leverages efficient spatio-temporal attention to model the complex dependencies underlying protein dynamics.
Through several key technical improvements, including historical context noise addition during training and inference, block-diffusion-style causal training, and architectural optimizations, STAR-MD achieves efficient, scalable training and stable trajectory generation, significantly improving over current state-of-the-art models.
We perform an exhaustive evaluation across multiple different simulation timescales ranging from to , providing a comprehensive assessment of conformation quality, coverage, and dynamic fidelity. Our analyses yield new insights into the limitations of current state-of-the-art models in long-horizon generation and offer valuable guidance for future model design.
2 Related Work
Protein Conformation Generation. Several models [jingAlphaFoldMeetsFlow2024, lewis2024scalableBioemu, wang2024proteinconfdiff] build on advances in protein structure prediction [jumper2021_af2, abramson2024accurate, lin2023evolutionary] and diffusion models [hoDenoisingDiffusionProbabilistic2020ddpm] to directly generate time-independent conformations. These methods provide an efficient alternative to MD by enabling parallel sampling. However, they capture only the equilibrium distribution of conformations and cannot model the temporal evolution of protein dynamics.
MD Trajectory Generation. To model the temporal evolution, operator-based methods [klein2024timewarp, costa2024equijump] aim to learn transport operators that predict conformations at lagged times. These methods approximate evolution as a Markovian process, and thus fail to capture the non-Markovian properties often present in partially observed systems such as protein dynamics data. In contrast, methods generating trajectories, akin to video generations, consider dependencies across multiple time frames. AlphaFolding [cheng20254d] incorporates higher-order information through additional “motion frames” and generates multiple future frames simultaneously. Similarly, MDGen [jing2024generative_mdgen] models the joint distribution of frames across 100-ns trajectories. However, both methods only capture the dependencies in a fixed context and prediction window. When generating longer trajectories through extension, they discard memory of earlier windows, breaking temporal consistency. To address this, ConfRover [shen2025simultaneous_confrover] adopts language model-style autoregression that can generate trajectories of arbitrary lengths while maintaining full memory via KV-caching. STAR-MD follows this approach to avoid the “memory break” in block-based models.
Scalable Structural and Temporal Modeling. Current structural models, including AlphaFolding and ConfRover, represent protein structures using single and pairwise features and process them with specialized architectures such as Pairformer and Invariant Point Attention (IPA) [jumper2021_af2]. While these modules are highly expressive and preserved the required roto-translational symmetries, they are computationally and memory intensive. This limits model efficiency and scalability, especially when extended to modeling trajectories across multiple time frames. MDGen sidesteps this limitation by anchoring trajectories to key frames, encoding only single embeddings with standard transformers. While this allows modeling of long trajectories (e.g., 250 frames for ATLAS), the reliance on key frames limited flexibility and model performance is suboptimal. Lastly, all the above approaches model dynamics between frames using interleaved spatial and temporal modules. While this decoupled design reduces computational cost, it limits the model’s expressive power to capture complex spatio-temporal relationships. In contrast, STAR-MD proposes to employ joint spatio-temporal attentions on the single embeddings’ direct space-time processing while keeping the memory footprint manageable.
3 Method
To generate long and physically plausible protein trajectories, we propose Spatio-Temporal Autoregressive Rollout for Molecular Dynamics (STAR-MD), a spatio-temporal -diffusion model that operates within an autoregressive framework. In the following sections, we first describe the overall autoregressive diffusion framework, then detail the architecture of our diffusion model, and finally explain the training and inference procedures.
3.1 Autoregressive SE(3)-Diffusion Models for Protein Trajectories
Autoregressive Trajectory Generation. We aim to generate protein conformation trajectories in an autoregressive fashion , where represents the time interval between frames. This formulation amortizes trajectory generation into a frame-level process conditioned on the entire past history, enabling the model to generate future dynamics of arbitrary length while maintaining a flexible memory representation learned from data. Causal transformers, widely used in language models for efficient autoregressive sequence modeling [touvron2023llama], provide a natural architectural choice for this task.
SE(3) Diffusion Models. To model the generation process of each frame , we use a (conditional) diffusion model on the Riemannian manifold , where which captures translation positions and rotations for each amino-acid residue in a protein sequence of length [yim2023se, wang2024proteinconfdiff, shen2025simultaneous_confrover]. The forward diffusion process independently corrupts translations and rotations across the diffusion time
where IGSO3 is the isotropic Gaussian on . Given a noisy structure , diffusion time , and condition , a denoising score network
is trained via denoising score matching to predict the noise added to both components. During inference, the learned score function can be used to sample from the reverse diffusion process and generate protein structures.
To capture temporal dependencies needed for autoregressive generation, the condition for generating the current frame should incorporate all past history . Specifically, we design an efficient and expressive causal diffusion transformer network to compute from the preceding clean frames and the current noisy frame . This design differs from prior works such as MDGen [jing2024generative_mdgen] & Alphafolding [cheng20254d] which do not include autoregressive conditioning, and ConfRover [shen2025simultaneous_confrover] which compresses all preceding frames into a static condition . In the following section, we detail the design of and how it efficiently integrates spatio-temporal information from preceding frames.
3.2 Architecture
In this section, we describe the architecture choices underlying STAR-MD. The main features of this architecture lie in the diffusion decoder, which uses spatio-temporal attention for autoregressive conditioning on previous frames and block diffusion allowing for efficient training and KV caching.
Input Module. We mirror the FrameEncoder module in shen2025simultaneous_confrover, starting with a frozen OpenFold FoldingModule for sequence-level single-residue and pairwise features . We incorporate these time-independent features with amino-acid metadata and pairwise distance within each frame to obtain time-dependent features , which are used as initial input to the diffusion blocks.
Diffusion Blocks. Our diffusion module contains submodules stacked into blocks:
1.
Invariant Point Attention layers [jumper2021_af2] update single features using pair features and noisy frames . This step is independent for each frame.
2.
Joint Spatio-Temporal Attention layers update the single-residue features by attending to those from previous frames , which allows for exchange of temporal informaiton.
3.
Pair features are updated from single features using the EdgeTransition layer from jumper2021_af2. Noisy frames are updated from single features using an MLP-based backbone update.
After stacking the above blocks, the final update of the coordinates is used as the score function prediction and fed to the denoising score matching loss.
Joint Spatio-Temporal Attention. Our joint spatio-temporal (ST) attention mechanism integrates information across the temporal dimensions, departing from previous models that employ factorized, “space-then-time” attention which imposes a restrictive bias that spatial and temporal dependencies are separable. Instead, our ST attention operates on tokens representing residue-frame pairs , allowing it to directly model non-separable relationships, such as how motion at one residue is coupled to past motion at a distant site. 2D Rotary Position Embedding (2D-RoPE) [heo2024rotary] is used to embed residue and frame indices, which allows extrapolation past the training number of frames.
This architectural choice is key to our model’s scalability. We analyze the per-layer computational complexity of our ST attention () against general ST (i.e., “space-then-time”) architectural paradigms (see Section˜9 for details). Most baseline architectures employ a Pairformer [jumper2021_af2] backbone, incurring a cubic spatial complexity due to expensive triangular attention operations. We consider two representative baseline configurations: (1) Pairformer + Pair Temporal Attention (e.g., ConfRover [shen2025simultaneous_confrover]): This configuration scales as . Here, the pairwise temporal attention alone incurs cost, matching that of STAR-MD, but the additional cubic spatial overhead makes it significantly more expensive. Furthermore, this approach incurs a quadratic memory cost for caching pairwise key-value states during autoregressive generation, which becomes a severe bottleneck for long rollouts. In contrast, STAR-MD only requires memory for caching single-residue features. (2) Pairformer + Single Temporal Attention: This configuration scales as . While the temporal scaling is theoretically superior to STAR-MD’s in the limit , the cubic spatial cost dominates in practical regimes. STAR-MD thus offers a favorable trade-off, enabling efficient modeling of long trajectories for large proteins without the cubic spatial overhead.
3.3 Additional Model Details
Block-causal Attention for Efficient Training. Traditional video diffusion models [peebles2023scalable, ma2024sit] denoise all frames simultaneously, using a fixed window size. For autoregressive generation with variable context length, the model denoises one frame at a time, attending to clean history contexts. This is straightforward during inference, where we cache clean frames for efficient autoregressive generation. However, given the sequential nature of this process, extra work is needed for parallel training. To simultaneously denoise all tokens with clean history context, we follow [arriola2025block, teng2025magi, deng2024causalfusion] to concatenate clean and noisy frames as input sequence and employ a special block-wise attention pattern: all frames only attend to clean history frames, preserving the causal structure during training. At the cost of doubling the sequence length, the model can optimize the training loss for all frames in a single forward pass, aligning parallel teacher-forcing training with sequential autoregressive inference.
Contextual Noise Perturbation for Robust Rollouts. To mitigate compounding errors during long autoregressive rollouts, we apply a contextual noise perturbation technique inspired by Diffusion Forcing [chen2024diffusion, song2025history]. During training, we perturb historical context frames by applying the forward diffusion process (Eqs. 1-2) with a small, randomly sampled noise level to obtain noisy contexts . The model then learns to predict frame conditioned on this perturbed history. Critically, at inference time, we apply the same noise perturbation: after generating frame , we add noise at level before using it as context for subsequent frames. This training-inference consistency ensures that the model experiences similar input distributions during deployment as during training, making it robust to its own prediction errors and preventing compounding drift in long rollouts.
Continuous-time Conditioning. To handle the vast range of timescales in protein motion, we draw the physical stride ns independently for every training example. By conditioning the network on through adaptive layernorm (AdaLN, [peebles2023scalable]), the model learns to modulate its internal activations as a function of both diffusion progress and physical timescale. Crucially, this training strategy decouples the physical duration of a trajectory from the number of frames in the context window. Even when restricted to small context windows during training (as most methods are), randomly sampling large strides exposes the model to relative time deltas spanning orders of magnitude. This allows the network to learn long-range temporal dependencies without the computational cost of training on long sequences or the need for complex context-length extrapolation techniques.
3.4 Theoretical Justification for STAR-MD Architecture
In this subsection, we use the Mori-Zwanzig formalism [mori1965transport, zwanzig1961memory] (See Section˜8 for more details) to provide theoretical justification for two key aspects of our architecture: the necessity of temporal history in coarse-grained modeling, and the specific requirement for joint spatio-temporal attention arising from the removal of explicit pairwise features.
Atomistic molecular dynamics simulations evolve according to Hamiltonian mechanics and are Markovian in the full phase space of atomic positions and momenta. However, practical generative models must operate on coarse-grained representations, such as per-residue coordinates for protein structures. The Mori-Zwanzig formalism shows that projecting Markovian dynamics onto coarse-grained variables yields a Generalized Langevin Equation with three terms (Equation˜5): a Markovian component, a memory kernel encoding history dependence, and a stochastic force representing eliminated degrees of freedom. Consequently, accurate modeling of coarse-grained protein dynamics requires non-Markovian models that incorporate temporal history to capture memory effects arising from eliminated degrees of freedom. Our work makes this connection explicit, providing the theoretical justification underlying recent temporal architectures in trajectory modeling [jing2024generative_mdgen, cheng20254d, shen2025simultaneous_confrover].
Existing models typically rely on explicit pairwise features to capture spatial structure, but this incurs prohibitive or costs (see Section˜3.2). STAR-MD circumvents this by projecting out pairwise features, a design choice we analyze through our Memory Inflation proposition (Section˜8.2.3, Section˜8). We show that removing explicit spatial correlations “inflates” the memory kernel for the remaining residues, necessitating a significantly richer temporal history to compensate. Crucially, this inflated kernel exhibits non-separable spatio-temporal coupling (Remark˜2), meaning spatial and temporal modes cannot be factorized. This theoretical insight directly motivates our architecture: instead of the interleaved spatial and temporal blocks used in prior work, STAR-MD employs joint spatio-temporal attention to approximate this complex, non-separable memory kernel, balancing physical fidelity with computational scalability.
4 Experiments
We conduct a comprehensive set of experiments to evaluate STAR-MD’s ability to generate long-horizon protein dynamics. First, we benchmark STAR-MD against state-of-the-art models on the standard trajectory generation task (Section 4.2). Next, we assess its extrapolation capabilities by generating trajectories over longer, microsecond-scale horizons not seen during training (Section 4.3). Finally, we perform a series of analyses and ablations to investigate model stability, scalability, and the contributions of our key design choices (Section 4.4).
4.1 Experimental Setup
Dataset. ATLAS dataset [vander2024atlas] contains MD trajectories for 1390 structurally diverse proteins. We use the standard train/val/test splits from prior works [jing2024generative_mdgen, shen2025simultaneous_confrover] to assess model performance in a transferable setting. To evaluate long-horizon generation–a key focus of our work–we extend the benchmark by running new MD simulations to produce 250 ns and trajectories for selected proteins, using the original ATLAS simulation protocols for consistency. Further details are provided in Appendix 10.
Implementations. We follow the training procedure described in Section 3.3 and generate trajectories using the configurations listed in Table 4. We compare STAR-MD with three state-of-the-art trajectory generative models trained on ATLAS: AlphaFolding [cheng20254d], MDGen [jing2024generative_mdgen], and ConfRover [shen2025simultaneous_confrover]. All models parametrize proteins using backbone rigids (translations and rotations) with torsional angles for side-chain atoms. To standardize evaluation, trajectories from variable-stride models (STAR-MD, ConfRover) are generated directly at the required intervals. Trajectories from fixed-stride models (AlphaFolding, MDGen) are first generated at their native resolution and then subsampled. Finally, we include an oracle reference based on MD simulations run independently to represent the target performance. All trajectories are aligned to the starting frame of the reference ATLAS trajectory via C superposition prior to analysis (Section˜11). Specific parameters for each benchmark are detailed in the relevant sections below, with full implementation details in Appendix 13.
Structural Quality: We assess the physical plausibility of generated conformations using a hierarchy of geometric and stereochemical metrics. First, we perform coarse-grained checks for C-C clashes (non-bonded atoms too close) and chain breaks (consecutive C atoms too far apart). Second, we use the MolProbity suite [chen2010molprobity] for fine-grained, all-atom analysis of backbone Ramachandran and side-chain rotamer outliers. We report three distinct validity metrics based on different criteria: C-level Validity, passing only C checks; All-Atom Validity, passing only MolProbity checks; and Combined Validity, passing all checks simultaneously. See Appendix 11 for details on the thresholds.
Conformational Coverage: To evaluate how well generated trajectories explore the conformational space of the reference MD simulation, we follow the protocol of shen2025simultaneous_confrover. We project all conformations into a low-dimensional space defined by the principal components of the reference trajectory and compute the Jensen-Shannon Divergence (JSD) and conformation recall between the distributions. To ensure that coverage reflects physically plausible exploration, we report these metrics computed exclusively on structurally valid conformations (“Cov Valid”).
Dynamic Fidelity: We assess temporal coherence using four metrics. First, we use tICA lag-time correlation to quantify the preservation of slow collective variables [molgedey1994separation, perez2013identification, shen2025simultaneous_confrover]; crucially, this is computed only on valid transitions to avoid artifacts from broken structures. Additionally, we evaluate RMSD to measure the magnitude of structural change over varying intervals, autocorrelation to assess temporal memory, and VAMP-2 score to evaluate how well slow dynamical modes of the system are captured. For last three metrics, we report the deviation from MD reference, with details in Appendix 11.
4.2 Standard Benchmark: Trajectory Generation
We first evaluate STAR-MD on the standard ATLAS benchmark. Following the protocol of shen2025simultaneous_confrover, we generate trajectories of 80 frames at a 1.2 ns interval. Results are summarized in Table˜1. It is important to note that due to its high computational cost, AlphaFolding failed to run on the four largest proteins in the test set, highlighting its scalability limitations even on this standard benchmark. On the full test set, STAR-MD, achieves a superior balance of conformational coverage, dynamic fidelity, and structural quality.
The results demonstrate clear performance deficiencies in all baseline models. Baselines like MDGen and ConfRover exhibit low conformational coverage (0.30 and 0.42 Recall, respectively) and produce a substantial number of structurally invalid frames, with validity rates of only 64.9% and 49.7%, respectively. AlphaFolding achieves a higher raw recall (0.51), but this metric is misleading as a negligible 0.11% of its generated frames are structurally valid; when controlled for valid structures, this recall drops to 0.01, rendering AlphaFolding as a very unreliable model for this task. In contrast, STAR-MD achieves the highest conformational coverage (0.58 recall) and structural validity (86.36%) of any generative model, significantly narrowing the performance gap to the ground-truth MD simulation which serves as the oracle for this task.
| Cov Valid | Dynamic Fidelity | Validity | |||||||
| Model | JSD | Rec | tICA | RMSD | AutoCor | VAMP-2 | CA% | AA% | CA+AA% |
| MD (Oracle) | 0.31 | 0.67 | 0.17 | 0.00 | 0.00 | 0.02 | 98.37 | 98.07 | 96.43 |
| MDGen | 0.560.01 | 0.280.01 | 0.120.00 | 0.38 0.01 | 0.05 0.00 | 0.38 0.01 | 71.831.90 | 95.030.59 | 68.312.20 |
| AlphaFolding111AlphaFolding results are evaluated on 78/82 proteins due to out-of-memory error for 4 large proteins. | 0.590.01 | 0.200.01 | N/A | 3.31 0.06 | 0.12 0.01 | 1.56 0.01 | 10.580.09 | 0.820.10 | 0.470.04 |
| ConfRover | 0.520.01 | 0.360.01 | 0.150.01 | 0.20 0.00 | 0.08 0.00 | 0.47 0.02 | 56.940.52 | 92.470.25 | 52.060.36 |
| STAR-MD | 0.430.01 | 0.540.01 | 0.170.00 | 0.07 0.02 | 0.02 0.00 | 0.10 0.02 | 86.810.64 | 98.180.05 | 85.290.62 |
In Figure 2(a) and 8, we compare the level of conformation changes at different lag times, as an indicator how well the model captures the ground truth motion changes. STAR-MD shows improved coherence with MD reference trajectories, while AlphaFolding significantly overestimates the motion and MDGen and ConfRover tend to underestimate the dynamics level.
Finally, Figure 2(b) provides a qualitative assessment of conformational coverage using the principal component projection for an example ATLAS protein. STAR-MD exhibits better significantly better exploration and diversity than comparison methods and is able to populate two distinct modes.
| Cov Valid | Dynamic Fidelity | Validity | ||||||
| Model | Time | JSD | Rec | RMSD | AutoCor | CA% | AA% | CA+AA% |
| 0.26 | 0.75 | 0.01 | 0.00 | 99.53 | 96.83 | 96.36 | ||
| MD (Oracle) | 0.23 | 0.91 | 0.00 | 0.00 | 96.25 | 86.50 | 82.75 | |
| MDGen | 0.520.01 | 0.380.01 | 0.48 0.01 | 0.25 0.01 | 63.252.10 | 87.831.13 | 56.602.10 | |
| 0.560.01 | 0.360.03 | 0.37 0.02 | 0.39 0.01 | 36.117.34 | 56.994.52 | 24.814.30 | ||
| Alphafolding | 0.570.03 | 0.200.03 | 1.760.05 | 0.140.01 | 8.960.25 | 0.940.12 | 0.630.16 | |
| 0.650.00 | 0.200.00 | 0.780.03 | 0.040.01 | 9.640.02 | 0.190.00 | 0.060.00 | ||
| ConfRover-W | 0.510.01 | 0.420.02 | 0.35 0.01 | 0.39 0.01 | 44.711.55 | 73.130.84 | 36.511.22 | |
| 0.550.02 | 0.450.02 | 0.33 0.02 | 0.38 0.03 | 54.741.79 | 62.323.43 | 36.911.39 | ||
| STAR-MD | 0.440.01 | 0.590.01 | 0.20 0.02 | 0.03 0.01 | 85.161.91 | 97.570.13 | 83.151.99 | |
| 0.460.01 | 0.610.02 | 0.13 0.02 | 0.10 0.02 | 88.471.09 | 89.810.65 | 79.931.04 | ||
4.3 Extrapolating for Long-Horizon Trajectories
The ability to generate stable, physically realistic trajectories over extended time horizons is critical for modeling functionally relevant protein dynamics. Most conformational transitions of biological interest occur beyond the timescale, where existing generative models begin to fail. We assess STAR-MD’s long-horizon capability through two increasingly challenging settings: simulation with 32 proteins and simulation with 8 proteins, with all proteins unseen during training. To rigorously evaluate temporal extrapolation, all models were trained exclusively on the ATLAS trajectories and received no fine-tuning for longer horizons. This setup directly tests the ability to generalize to dynamics far beyond the training data distribution. Since the ATLAS benchmark only provides trajectories up to , we generated our own molecular dynamics simulations for longer timescales to create proper reference data for evaluation. These extended simulations follow the same simulation protocols as ATLAS but continue the dynamics to longer timescales.
Due to its scaling issues, ConfRover could not be evaluated on either or generation tasks. This is because ConfRover performs temporal attention over pair features, requiring previous frames’ pair features to be stored in memory as KV cache (see Appendix 9 for more detailed analysis). As such, even with CPU offloading of the KV cache, ConfRover’s memory requirements exceeded our hardware limits (1869 GB CPU RAM, 8 H100 GPUs with 80GB VRAM each). As a result and for fair comparison, we utilize a variant of ConfRover with windowed attention with attention sink tokens (as described in xiao2023efficient) to reduce memory usage. We report results for this windowed variant, labeled ConfRover-W, in Table˜2.
Table˜2 summarizes the results for these two extended timescales. Remarkably, STAR-MD maintains stable and competitive performance at both the timescale and the challenging timescale, demonstrating the effectiveness of our joint spatio-temporal attention and contextual noise techniques for long-horizon stability. Further, STAR-MD exhibits controlled error accumulation, with low clash and break rates that remain consistent throughout the extended trajectory (Figure 3 (right), Section˜4.4). Among the methods achieving reasonable structural quality, STAR-MD attains the best balance of coverage and fidelity, nearing the oracle MD performance. In both the and case, AlphaFolding produces physically implausible, high-error structures despite good coverage metrics, while ConfRover-W and MDGen degrade significantly over the microsecond horizon. We further report kinetic metrics for the long trajectory cases in Figure˜8, where STAR-MD shows the best alignment with the RMSD and autocorrelation of the MD reference from among baseline methods.
4.4 Additional Analysis
Error Accumulation. A key challenge in long-horizon generation is error accumulation, where minor inaccuracies compound over time, leading to trajectory degradation. Figure 3 plots the average validity percentage over simulation time for the , , and simulations. As the simulation horizon extends, baseline models exhibit rapidly deteriorating performance, with structural validity declining sharply. In contrast, STAR-MD remains stable, maintaining high structural validity close to the MD oracle across all timescales. This demonstrates our model’s robustness against error accumulation, a quality largely attributable to the historical-context noise mechanism, which enables stable long-horizon autoregressive generation.
Varying temporal resolution for long-horizon generation. The generation task above uses a stride of 2.5 ns, resulting in 400 frames. Thanks to our continuous-time conditioning, we can generate at different temporal resolutions without retraining. We test this by generating a trajectory with a 1.2 ns stride, nearly doubling the number of frames to 850. Figure 5 shows that STAR-MD remains stable even with this much longer sequence of frames, maintaining high structural quality. This further underscores our model’s robustness to different sampling intervals and trajectory lengths, showing strong capability for stable, long-horizon autoregressive generation.
KV Cache Analysis. A critical limitation of ConfRover is the need to maintain the KV-cache of single and pair embedding for temporal attentions. In contrast, STAR-MD only requires maintaining the KV cache for singles. By only maintaining attentions among single representations, STAR-MD has a magnitudes lower memory footprints (see Figure˜5 for a comparison). This advantage allows our model to maintain a full KV caches in GPU memory without resorting to CPU offloading (e.g, additional overhead) or sliding window style KV caches (compromises temporal history).
4.5 Ablation Studies on Key Components
| Cov Valid | Dynamic Fidelity | Validity | |||||||
| Model | JSD | Rec | tICA | RMSD | AutoCor | VAMP-2 | CA% | AA% | CA+AA% |
| STAR-MD | 0.42 | 0.57 | 0.17 | 0.09 | 0.02 | 0.12 | 86.62 | 98.28 | 85.18 |
| w/o Noise | 0.43 | 0.54 | 0.18 | 0.04 | 0.11 | 1.22 | 77.82 | 97.76 | 76.12 |
| w/ Sep Attn | 0.46 | 0.46 | 0.17 | 0.09 | 0.05 | 0.49 | 87.95 | 98.34 | 86.70 |
| w/ Preproc Attn | 0.47 | 0.48 | 0.17 | 0.09 | 0.05 | 0.54 | 84.55 | 97.81 | 82.56 |
To validate our key design choices, we conduct systematic ablations of the main components of STAR-MD. Table 3 summarizes the quantitative results on the benchmark. Section˜15.2 contains full ablation tables for the , , and benchmarks.
ST attention vs. separable attention. We compare our joint ST attention with a separable alternative that processes spatial and temporal dimensions sequentially. The separable model suffers a significant drop in conformational coverage and a modest drop in dynamic fidelity. This confirms that joint ST attention is crucial for capturing the complex, non-separable spatio-temporal dependencies inherent in protein dynamics, which directly impacts the model’s ability to explore the correct conformational space.
ST layer placement. We evaluated placing the spatio-temporal attention module outside the diffusion decoder as a static conditioner which compresses the full trajectory history into a single conditioning vector, mirroring shen2025simultaneous_confrover. This variant underperforms our full model in both coverage and fidelity, confirming that integrating spatio-temporal attention directly within the diffusion module is essential for effective context utilization.
Historical-context noise. Removing the contextual noise perturbation leads to a substantial drop in structural quality. This highlights the importance of this technique for maintaining generation stability, a finding that is further supported by our long-horizon experiments in Section 4.3. Figure˜12 compares the CA+AA validity over time for models with and without contextual noise on the and benchmarks, showing that the noise helps maintain structural quality over long horizons.
These ablation studies demonstrate that each component of STAR-MD addresses a specific challenge in protein dynamics modeling. The ST attention provides the expressivity needed for complex spatio-temporal dependencies. The historical-context noise ensures stable long-horizon generation by preventing error accumulation.
5 Conclusion
We introduced STAR-MD, a novel causal diffusion transformer model for generating long-horizon protein dynamics trajectories. STAR-MD addresses the challenges of long-horizon generation through several key innovations: joint spatio-temporal attention that efficiently models complex space-time couplings, continuous-time conditioning for generation at arbitrary timescales, and a noisy-context training scheme that mitigates error accumulation. Our comprehensive evaluations demonstrate that STAR-MD achieves new state-of-the-art performance on the standard ATLAS benchmark, outperforming previous methods in conformational coverage, structural quality, and kinetic fidelity. More importantly, our model successfully extrapolates to generate stable, high-fidelity trajectories up to the microsecond regime, where prior models often fail due to computational costs or compounding errors. By balancing expressiveness and scalability, STAR-MD establishes an efficient and empirically strong foundation for modeling protein dynamics at biologically relevant scales, paving the way for accelerated exploration of complex biological processes.
Limitations and future work. While STAR-MD represents a significant step forward, there are avenues for future improvement. The temporal consistency of the generated trajectories, while strong, does not yet perfectly match that of oracle MD simulations. The model’s performance could be further enhanced by training on larger and more diverse MD simulation datasets, such as MDCATH [mirarchi2024mdcath]. Additionally, a promising direction for future work is to extend the model’s capabilities to simulate the dynamics of protein complexes or their interactions with small molecules, which are crucial for understanding biological processes and drug design.
6 Author Contributions
Y.S. and Q.G. conceived the study, designed the experimental protocols, and supervised methodology development. N.S., Y.L., and Y.S. developed the spatiotemporal causal diffusion architecture and analyzed model performance. P.L. and Y.S. conceptualized the contextual noise perturbation, which N.S. implemented. N.S., P.L. and R.B. conducted the theoretical verification of the Mori-Zwanzig formalism and validated the mathematical frameworks. All authors participated in the manuscript preparation and critical review for publication.
7 Acknowledgments
We would like to thank Chan Lu, Cheng-Yen (Wesley) Hsieh, Yi Zhou, Zaixiang Zheng, Yilai Li and other members of ByteDance for their valuable feedback and suggestions that helped improve this work.